# Identifying Fairness Issues in Automatically Generated Testing Content

### v1

@@ Abstract
Natural language generation tools are powerful and effective for generating content. However, language models are known to display bias and fairness issues, making them impractical to deploy for many use cases. We here focus on how fairness issues impact automatically generated test content, which can have stringent requirements to ensure the test measures only what it was intended to measure. Specifically, we identify test content that is focused on particular domains and experiences that only reflect a certain demographic or that are potentially emotionally upsetting; both of which could inadvertently impact a test-taker’s score. This kind of content doesn’t reflect typical biases out of context, making it challenging even for modern models that contain safeguards. We build a dataset of 621 generated texts annotated for fairness and explore a variety of methods for classification: fine-tuning, topic-based classification, and prompting, including few-shot and self-correcting prompts. We find that combining prompt self-correction and few-shot learning performs best, yielding an F1 score of .791 on our held-out test set, while much smaller BERT- and topic-based models have competitive performance on out-of-domain data. [^1] Code and dataset will be released upon publication.
@@ 1 Introduction
Figure 1: In (A), the generated question and expected answer unfairly advantage those with knowledge of Taylor Swift, while (B) more neutral. Our goal is to identify and filter content like (A) to ensure fair testing.
@@
Large language models (LLMs) have become ubiquitous in the space of natural language generation (NLG) due to recent advances in model capability Minaee et al. (2024). However, these improvements come with the potential for various negative societal impacts. These negative impacts include the generation of misinformation/propaganda, allocation harms of systems providing benefits only to certain groups of people, and representational harms revolving around bias and stereotyping. Natural language processing (NLP) models including LLMs are known to reflect and repeat harmful biases and stereotypes Hosseini et al. (2023); Bender et al. (2021); Hovy and Prabhumoye (2021); Nadeem et al. (2021), and research into how the community addresses the societal harms engendered by NLP technology is critical Wang et al. (2024); Dev et al. (2022); Blodgett et al. (2020). These issues are particularly dangerous when applied to language learning and assessment; tests with inherent biases have the potential to directly harm test takers by unfairly disadvantaging certain subgroups.
@@
Many of these types of bias in language generation are well studied. Biases based on gender Nemani et al. (2024); Devinney et al. (2022); Strengers et al. (2020); Wan et al. (2023), race Das and Balke (2022); Field et al. (2021), nationality Venkit et al. (2023), and disability Venkit et al. (2022) have been identified in language models, and many modern LLMs incorporate deliberate safeguarding measures in an attempt to alleviate these issues OpenAI et al. (2023); Anil et al. (2023).
@@
In the area of language assessment, there exists a tangential set of issues regarding fairness to test takers and score users Educational Testing Service (2022). Content that is irrelevant to the skills and abilities being tested should be avoided (Figure 1). This includes content that could disadvantage anyone based on their culture, location, or experiences (e.g., focusing on barbeques on the 4th of July could disadvantage test-takers who are unfamiliar with U.S. culture); their emotions (e.g., health hazards and diseases can evoke negative emotional responses among some people); their worldviews (e.g., luxury cruises or designer clothing may make some people feel excluded); and other issues. We refer to these types of issues as fairness issues.
Knowing how to better understand, detect, and mitigate bias related to fairness in NLG not only raises awareness of the issue but also enables researchers and developers to create more fair and inclusive NLP systems, evaluation metrics, and datasets in the language assessment space.
@@
Our goal is to build a system for identifying fairness-violating content in automatically generated texts. It is of course still necessary to have human review of content before it is publicly used, but by adding a filtering process after generation and before manual review, we can significantly reduce the time taken for reviewing and the chance that fairness-related content is mistakenly allowed. To accomplish this goal, we explore four different approaches: fine-tuning, topic-based classification, few-shot prompting, and prompt-self correction.
@@
Our methods need to adapt to new contexts: our definition of fairness is operationally defined by the particular testing context, and may not apply to others, so the guidelines, prompts, and models may not apply generally to new contexts. For this reason, we assess our methods on two held-out test sets and analyze how our methods could be applied to new contexts. We release our resulting dataset, consisting of 620 samples of which 19.4% contain fairness issues [^2] Each sample we used was rejected for deployment in actual tests. Using rejected samples for our experiments allows us to release the dataset: accepted stimuli cannot be made public., to facilitate improvements in the fairness-detection community.
@@
    We build and release a dataset of 620 generated samples for use in evaluating fairness detection methods.
@@
    We provide a new mechanism for prompting self-correction which yields significant improvements over other prompting strategies.
We start with data collection and analysis. We collect 620 samples over 7 different types of content generated using LLM prompting. We annotate each sample and identify whether it contains a fairness issue, and if so, whether that fairness issue pertains to knowledge, skill, or expertise or emotion (more on these categories and how they relate to fairness in Section 3). We then use this dataset to experiment with a series of classification models for identifying fairness issues.
We show that fine-tuning and filtering by topic can be cheap and effective options, although overall prompting strategies with GPT4 tend to be more effective. Few-shot prompting along with self-correcting prompt strategies yield strong performance with relatively little data, and combining both yields the best results on our in-domain test set, with an F1 score of .773. Interestingly, using a shorter, more generic prompt combined with our self-correction method yields the best result on our out-of-domain test set, with an F1 score of .462.
@@ 3 Data
  • Question: Are you planning to go to one of The Eras Tour shows?
  • Response: Yes—I love Taylor Swift!
If the task were to identify whether the response is an appropriate response to the question, even some native English speakers would likely get it wrong. This is because, in addition to needing to know features of English proficiency (in this case, the ability to infer gist, purpose, and basic context based on information stated in short spoken texts), one would also need to know about Taylor Swift and her concert tour. Thus, those familiar with Taylor Swift would have an unfair advantage in identifying the correct answer, and those who are unfamiliar with Taylor Swift would be unfairly disadvantaged.
@@
  • Question: Are you planning to go to the music concert on Friday?
  • Response: Yes—I’m looking forward to the performance!
@@ 3.1 Dataset
Our goal is to identify and mitigate these fairness issues in testing content. We build a dataset spanning seven different item or task types from standardized English language proficiency tests all generated using GPT4 OpenAI et al. (2023). Item and task types can contain up to four components: the stimulus (main text the question is based around), stem (question asked about the stimulus), key (correct answer to the stem), and distractors (a set of alternative answers that are not correct).
Fairness issues are possible in all components, but we focus on only the stimuli: these are the longest, most feature-rich components of the test content, and thus are most likely to reflect fairness and bias issues. Issues in the stimuli can leak through to stems, keys, and distractors, making the stimulus the likely culprit for the majority of issues.
@@
For each stimulus, we aim to identify whether or not the stimulus contains fairness/bias issues, and if so, what type of issue is present. We start with a dataset of automatically generated stimuli. These have been reviewed by experts, and each was rejected for use in live assessments. These rejected stimuli have the relevant language and structure: our goal is to identify which of those stimuli were rejected (at least in part) for fairness reasons. We employ content development experts to annotate these samples, yielding a binary classification between non-fairness and fairness-related rejections.
However, there are many different ways for bias and fairness considerations to impact individual stimuli. To better understand and mitigate these issues, we split them into two main categories:
@@
We annotate each sample flagged with a fairness issue with one or both of these categories. This allows further analysis to address these specific fairness categories and to better understand the impact of specific fairness issues.
Our dataset is comprised of stimuli from seven different item and task types: a summary of the collected data is shown in Table 1, with examples for each type in Appendix A. These stimuli represent various structures, depending on the item/task type: Read a Text Aloud, Talks, Announcements, and Text Completion stimuli are short text paragraphs, while Conversation stimuli involve turns between two or more speakers. Respond to Questions Using Information Provided and Respond to a Written Request task stimuli are structured content: the generation process creates text that is filled into a structured template; we use only the raw text.
Overall we collect 620 samples, of which 19.4% exhibit evidence of fairness issues, with 9.3% reflecting KSA issues and 12.2% Emotion issues.
We build a validation set of 48 samples reflecting a balance of the item and task types from the training types (Read a Text Aloud, Talks, Text Completion, and Respond to Questions Using Information Provided), and an equal-sized "in-domain" dataset from these stimuli is held separately for testing. These datasets contain an even number of positive and negative classes for fairness evaluations: as our goal is to able to identify positive cases where fairness issues exist, we intend for our validation and test sets to have a substantial number of this class. We use the three remaining types (Conversations, Respond to a Written Request, and Announcements) as a separate "out-of-domain" test set to evaluate performance on unseen content.
4 Methods
Our goal is to detect whether a generated stimulus contains an issue as a binary classification task. We experiment with standard transformer-based classification baselines, topic detection, and a variety of GPT4-based prompting, including methods for automatic prompt-self correction. We describe each method below: each is tuned on the validation set, and we report the best model performance on that set. We then evaluate model performance on two separate test sets in Section 5.
@@ 4 Methods
As a baseline, we fine-tune standard pre-trained transformer models for sequence classification. We experiment with bert-base-cased, bert-large-cased, and roberta-base. We perform a hyperparameter search and report results using the model with the best-performing results.
@@
We first identify topics found within the data. We use BERTopic Grootendorst (2022), a state-of-the-art neural topic modeling framework to extract topic representations from two sources of training data: (1) all samples from the training partition of our dataset and (2) our fairness guidelines. In this method, SentBERT Reimers and Gurevych (2019) converts each training document into a dense vector representation which are grouped by semantic similarity, creating clusters that represent different topics. For each of the two training sets, topic descriptions made up of the most important words in a cluster are generated for the clusters containing at least five supporting documents. We manually assess each topic description for themes that should be avoided based on their relation to known fairness issues and which topics are acceptable. Finally, for each unseen sample in test and validation datasets, we make predictions based on the single nearest topic cluster. If a sample falls within the boundaries of restricted topics, it is classified as a violation.
@@
bert-base-cased 1.0 .292 .452
bert-large-cased .923 .500 .649
roberta-base .923 .500 .649
@@
Topic-data 0.786 .458 0.579
Topic-guidelines 1.0 .042 0.08
@@
Results for these methods are shown in Table 2. The fine-tuned bert-based models perform fairly well, with F1 scores for bert-large-cased and roberta-base both nearing .65. The Topic-based Filtering models are worse, with the data-based system yielding an F1 score of .579. In all cases precision is much higher than recall: these models are conservative with predictions.
@@ 4.1 Prompting
We initially experiment with five different “base” prompts. We pair these with stimuli and use GPT4 to return “True” if the stimulus contains a fairness issue and “False” otherwise. These prompts represent different strategies [^3] Prompts in Appendix B:
  • generic (short)  53 tokens: Drawing from general knowledge of fairness and bias in LLMs, we write a generic prompt designed to combat attested LLM biases. This prompt is designed as weak baseline. Our goal is to determine if short, simple prompt can capture relevant issues, and whether or not it can be easily improved via self-correction or few-shot learning (Sections 4.1 and 4.1)
@@
  • guideline (short)  197 tokens: We craft a prompt based on guidelines for writing fair assessments. Using documentation that defines what constitutes fair assessment items and how to write them, we build a prompt capturing the important components of a fair question based on guidelines. The goal of this prompt is to determine if human-written guidelines based on theoretical issues will accurately capture these issues in real data.
@@
  • data-driven  142 tokens: We craft a prompt based on annotations in our data. We identify what topics and language cause fairness issues in the training data and build the prompt to reflect both these issues specifically, and how they might generalize to unseen item/task types and topics. This method is hypothesized to be the most effective, as it will address known issues in the data, but may not extend to unseen data, as it is built specifically around the given training samples.
@@
Initial experiments on the validation set showed two things: the generic (long) prompt performs similarly to the generic (short) in all cases, and the guideline (long) prompt failed to identify any fairness issues. We therefore focus our efforts on the three other prompts: generic (short)guideline (short)  and data-driven.
A primary issue in identifying the fairness issues we are concerned with is that they are only applicable in the context of certain testing environments. The guidelines and data we use are specific to a certain test: for other tests, other environments, and other contexts, different definitions of fairness and different alignments will be applicable. For this reason, the ability to adapt to new environments is critical in applying fairness constraints. We aim to build a system that can learn to adapt to new guidelines with minimal information.
@@
We leverage our annotated data using a prompt self-correction strategy. We start with a base prompt and iterate through a batch of samples taken from our training data. For each sample, we insert text into the prompt, perform a pass through GPT4, and evaluate the resulting output. If it’s correct, we continue. If not, we instruct the LLM to make adjustments to the prompt, either to disallow the sample (when the error is a false negative) or to allow it (when the error is a false positive). This process is run up to $e$ epochs, or stopped early if accuracy reaches 1 or the predictions don’t change. We run over $b$ batches of $n$ samples randomly drawn from the training data. We use the best-scoring prompt from the final batch for evaluation. [^4] For an example of the process, see Appendix C
@@
We run each of our experiments on the validation set to tune the prompts and parameters to optimize F1 score. Figure 2 shows the best results on the validation set. We explore each model’s effectiveness on unseen data in Section 5.
Figure 2: F1 scores on the validation set for each prompting method. Note that for generic (short)  the F1 score was 0. Full results in Appendix D
The base generic prompt fails, as the traditional bias and stereotyping issues are less likely to occur in our generated content, and the fairness issues we are concerned with are unlikely to be deemed as problematic out of context. Using a simplified version of our guidelines yields .364 F1 score for identifying fairness issues. The data-driven based on observations in the training data yields much better results (.7 F1). However, this may not extend well to novel cases, as the prompt is driven purely by our validation data.
Few-shot learning displays some interesting properties: we see significant improvements across all three prompts, using 3 samples (this yielded the best results across all validation runs). Even the minimal generic (short) prompt rises to over .6 F1 with minimal few-shot prompting.

### v2

@@ Abstract
Natural language generation tools are powerful and effective for generating content. However, language models are known to display bias and fairness issues, making them impractical to deploy for many use cases. We here focus on how fairness issues impact automatically generated test content, which can have stringent requirements to ensure the test measures only what it was intended to measure. Specifically, we review test content generated for a large-scale standardized English proficiency test with the goal of identifying content that only pertains to a certain subset of the test population as well as content that has the potential to be upsetting or distracting to some test takers. Issues like these could inadvertently impact a test taker’s score and thus should be avoided. This kind of content does not reflect the more commonly-acknowledged biases, making it challenging even for modern models that contain safeguards. We build a dataset of 601 generated texts annotated for fairness and explore a variety of methods for classification: fine-tuning, topic-based classification, and prompting, including few-shot and self-correcting prompts. We find that combining prompt self-correction and few-shot learning performs best, yielding an F1 score of 0.79 on our held-out test set, while much smaller BERT- and topic-based models have competitive performance on out-of-domain data. [^1] Code and dataset available at https://github.com/EducationalTestingService/fairness-detection.
@@ 1 Introduction
Figure 1: In (A), the generated question requires knowledge of what The Eras Tour is to identify the correct answer. Even native English speakers would likely not be able to identify the correct response if they were not familiar with Taylor Swift. In (B), the generated question does not require specific background knowledge, so test takers would not need to use specialized knowledge to identify the correct answer. Our goal is to identify and filter content like (A) to help ensure fair testing.
@@
Large language models (LLMs) have become ubiquitous in the space of natural language generation (NLG) due to recent advances in model capability Minaee et al. (2024). However, these improvements come with the potential for various negative societal impacts. These negative impacts include the generation of misinformation/propaganda, allocation harms of systems providing benefits only to certain groups of people, and representational harms revolving around bias and stereotyping. Natural language processing (NLP) models–including LLMs–are known to reflect and repeat harmful biases and stereotypes Hosseini et al. (2023); Bender et al. (2021); Hovy and Prabhumoye (2021); Nadeem et al. (2021), and research into how the community addresses the societal harms engendered by NLP technology is critical Wang et al. (2024); Dev et al. (2022); Blodgett et al. (2020).
@@
Many of these types of bias in language generation are well-studied. Biases based on gender Nemani et al. (2024); Devinney et al. (2022); Strengers et al. (2020); Wan et al. (2023), race Das and Balke (2022); Field et al. (2021), nationality Venkit et al. (2023), and disability Venkit et al. (2022) have been identified in language models, and many modern LLMs incorporate deliberate safeguarding measures in an attempt to alleviate these issues OpenAI et al. (2023); Anil et al. (2023).
@@
In the area of language assessment, there exists a tangential set of issues regarding fairness to test takers and score users Educational Testing Service (2022). These issues are particularly dangerous when applied to language learning and assessment; tests with inherent biases have the potential to compromise the validity of the test. Therefore, content that is irrelevant to the skills and abilities the test is intended to measure should be avoided (Figure 1). This includes content that could disadvantage anyone based on their culture, location, or experiences (e.g., focusing on barbeques on the 4th of July could disadvantage test-takers who are unfamiliar with U.S. culture); their emotions (e.g., health hazards and diseases can evoke negative emotional responses among some people); their worldviews (e.g., luxury cruises or designer clothing may make some people feel excluded); and other factors. We refer to these types of issues as fairness issues. Knowing how to better understand, detect, and mitigate bias related to fairness in NLG not only raises awareness of the issue but also enables researchers and developers to create more fair and inclusive NLP systems, evaluation metrics, and datasets in the language assessment space.
@@
Our goal is to build a system for identifying fairness-violating content in automatically generated texts. It is of course still necessary to have human review and revision of the content, but by adding a filtering process after generation and before manual review, we can significantly reduce the time taken for reviewing and the chance that fairness-related content is mistakenly allowed. To accomplish this goal, we explore four different approaches: fine-tuning, topic-based classification, few-shot prompting, and prompt-self correction.
@@
Our methods need to adapt to new contexts: our definition of fairness is operationally defined by the particular testing context, and may not apply to others, so the guidelines, prompts, and models may not apply generally to new contexts. For this reason, we assess our methods on two held-out test sets and analyze how our methods could be applied to new contexts. We release our resulting dataset, consisting of 620 samples, of which 19.4% contain fairness issues [^2] Each sample we used was rejected for deployment in actual tests. Using rejected samples for our experiments allows us to release the dataset: accepted stimuli cannot be made public., to facilitate improvements in the fairness-detection community.
@@
    We release a dataset of 601 samples for use in evaluating fairness detection methods.
@@
    We provide a new mechanism for prompting self-correction, which yields significant improvements over other prompting strategies.
We start with data collection and analysis. We collect 620 samples over seven different types of content generated using LLM prompting. We annotate each sample and assess whether it contains a fairness issue, and if it does, whether that fairness issue pertains to knowledge, skill, or expertise or emotion (more on these categories and how they relate to fairness in Section 3). We then use this dataset to experiment with a series of models for classifying fairness issues.
We show that fine-tuning and filtering by topic can be cheap and effective options, although prompting strategies with GPT4 tend to be more effective. Few-shot prompting along with self-correcting prompt strategies yield strong performance with relatively little data, and combining both yields the best results on our in-domain test set, with an F1 score of .773. Interestingly, using a shorter, more generic prompt combined with our self-correction method yields the best result on our out-of-domain test set, with an F1 score of .462.
@@ 3 Problem Motivation
  • Question: You went to one of The Eras Tour shows, didn’t you?
  • Response: Yes–I love Taylor Swift!
If the task were to identify whether the response is an appropriate response to the question, even some native English speakers would likely get it wrong. This is because, in addition to needing to know features of English proficiency (in this case, the ability to infer gist, purpose, and basic context based on information stated in short spoken texts), one would also need to know about Taylor Swift and her concert tour. Thus, those familiar with Taylor Swift would have an unfair advantage in identifying the correct answer.
@@
  • Question: You went to the music concert, didn’t you?
  • Response: Yes–it was a great performance!
@@
4 Methods
Our goal is to detect whether a generated stimulus contains an issue as a binary classification task. We build a dataset of texts labeled for potential fairness issues and explore potential detection methods.
@@ 4.1 Dataset
Our goal is to identify and mitigate these fairness issues in testing content. We build a dataset spanning seven different item or task types from standardized English language proficiency tests all generated using GPT4 OpenAI et al. (2023). Item and task types can contain up to four components: the stimulus (main text the question is based on), stem (question asked about the stimulus), key (the correct answer to the stem), and distractors (a set of alternative answers that are incorrect).
Fairness issues are possible in all components, but we focus on only the stimuli, which are typically the longest, most feature-rich components of the test content, and thus are most likely to reflect fairness and bias issues. Issues in the stimuli can leak through to other components, making the stimulus the source of the majority of fairness issues.
@@
For each stimulus, we aim to identify whether or not the stimulus contains fairness/bias issues, and if so, what type of issue is present. We start with a dataset of automatically generated stimuli. These stimuli were generated using prompting and different versions of GPT: the prompts were iteratively improved with the goal of improving the overall quality of the stimuli. During this process, each stimulus was evaluated by the test’s content development experts. For this work, the stimuli used were rejected by the reviewers, allowing us to provide them publicly and explore their use for fairness detection. These rejected stimuli typically have the relevant language and structure, so our goal is to identify which of those stimuli were rejected (at least in part) for fairness reasons. We employ content development experts to annotate these samples, yielding a binary classification between non-fairness and fairness-related rejections.
However, there are different ways for bias and fairness considerations to impact individual stimuli. To better understand and mitigate these issues, we separated them into two main categories:
@@
Each sample that is flagged for fairness is annotated for one or both of these categories. This allows further analysis to address these specific fairness categories and to better understand the impact of specific fairness issues.
Our dataset is comprised of stimuli from seven different item and task types: a summary of the collected data is shown in Table 1, with examples for each type in Appendix A. These stimuli represent various structures, depending on the item/task type: Read a Text Aloud, Talks, and Text Completion stimuli are short text paragraphs, while Conversation stimuli involve turns between two or more speakers. Respond to Questions Using Information Provided and Respond to a Written Request task stimuli are structured content: the generation process creates text that is filled into a structured template; we use only the raw text.
Overall we collect 601 samples, of which 19.6% exhibit evidence of fairness issues, with 9.5% reflecting KSA issues and 12.3% Emotion issues. We build a validation set of 48 samples reflecting a balance of the item and task types from the training types (Read a Text Aloud, Talks, Text Completion, and Respond to Questions Using Information Provided), and an equal-sized "in-domain" dataset from these stimuli is held separately for testing. These datasets contain an even number of positive and negative classes for fairness evaluations. As our goal is to be able to identify positive cases where fairness issues exist, we intend for our validation and test sets to have a substantial number of this class. We use the two remaining types (Conversations, Respond to a Written Request) as a separate "out-of-domain" test set to evaluate performance on unseen content.
4.2 Experiments
We experiment with standard transformer-based classification baselines, topic detection, and a variety of GPT4-based prompting, including methods for automatic prompt-self correction. We describe each method below: each is tuned on the validation set, and we report the best model performance on that set. We then evaluate model performance on two separate test sets in Section 5.
@@ 4.2 Experiments
We fine-tune standard pre-trained transformer models for sequence classification. We experiment with bert-base-cased, bert-large-cased Devlin et al. (2019), roberta-base, Liu et al. (2019) and deberta-base He et al. (2021) models. We perform a hyperparameter search on our validation set for each model, finding that a learning rate of 2e-5 over 2-4 epochs generally performs best, and report results using the model with the best performance.
@@
We first identify topics found within the data. We use the topic modeling framework BERTopic Grootendorst (2022) to extract topic representations from two sources of training data: (1) all samples from the training partition of our dataset and (2) our fairness guidelines. In this method, SentBERT Reimers and Gurevych (2019) converts each training document into a dense vector representation which are then grouped by semantic similarity, creating clusters that represent different topics. For each of the two training sets, topic descriptions made up of the most important words in a cluster are generated for the clusters containing at least five supporting documents. We manually assess each topic description for themes that should be avoided based on their relation to known fairness issues and which topics are acceptable. Finally, for each unseen sample in test and validation datasets, we make predictions based on the single nearest topic cluster. If a sample falls within the boundaries of restricted topics, it is classified as a violation.
@@
bert-base-cased 1.00 0.29 0.45
bert-large-cased 0.92 0.50 0.65
roberta-base 0.92 0.50 0.65
deberta-base 1.00 0.63 0.77
@@
Topic-data 0.79 0.46 0.58
Topic-guidelines 1.00 0.04 0.10
@@
Results for these methods are shown in Table 2. The fine-tuned bert-based models perform fairly well, with F1 scores for bert-large-cased and roberta-base both around 0.65, and deberta-base showing exceptional performance with an F1 score of 0.77. The Topic-Based Filtering models are worse, with the data-based system yielding an F1 score of 0.58. In all cases, precision is much higher than recall; these models are conservative with predictions.
@@ 4.3 Prompting
We initially experiment with five different “base” prompts. We pair these with stimuli and use GPT4 to return “True” if the stimulus contains a fairness issue and “False” otherwise. These prompts represent different strategies [^3] Prompts in Appendix B.:
  • generic (short)  53 tokens: Drawing from general knowledge of fairness and bias in LLMs, we write a generic prompt designed to combat attested LLM biases. This prompt is designed as a weak baseline. Our goal is to determine if a short, simple prompt can capture relevant issues, and whether or not it can be easily improved via self-correction or few-shot learning (Sections 4.3 and 4.3)
@@
  • guideline (short)  197 tokens: We craft a prompt based on guidelines for writing fair assessments. Using documentation that defines what constitutes fair assessment items and how to write them, we build a prompt capturing the important components of a fair question. The goal of this prompt is to determine whether human-written guidelines based on theoretical issues will accurately capture these issues in real data.
@@
  • data-driven  142 tokens: We craft a prompt based on annotations in our data. We identify which topics and language cause fairness issues and build the prompt to reflect how they might generalize to unseen item/task types and topics. This method is hypothesized to be the most effective, as it will address known issues in the data but may not extend to unseen data, as it is built specifically around the given training samples.
@@
Initial experiments on the validation set revealed two insights: the generic (long) prompt performs similarly to the generic (short) in all cases, and the guideline (long) prompt is ineffective. We therefore focus our efforts on the three other prompts: generic (short)guideline (short)  and data-driven.
A primary issue in identifying the fairness issues we are concerned with is that they are only applicable in the context of certain testing environments. The guidelines and data we use are specific to a certain test; for other tests, other environments, and other contexts, different definitions of fairness and different alignments will be applicable. For this reason, the ability to adapt to new environments is critical in applying fairness constraints. We aim to build a system that can learn to adapt to new guidelines with minimal information.
@@
We leverage our annotated data using a prompt self-correction strategy (Figure 2). We start with a base prompt and iterate through a batch of samples from our training data. For each sample, we insert text into the prompt, perform a pass through GPT4, and evaluate the resulting output. If it is correct, we continue. If not, we instruct the LLM to make adjustments to the prompt, either to disallow the sample (when the error is a false negative) or to allow it (when the error is a false positive), using the following prompts:
  • Positive sample incorrectly classified as negative: The previous text was classified as not having any fairness issues, returning False. However, it did have an issue. Think about what topics or issues in the text might be unfair. Improve the original fairness instructions by adding or modifying restrictions so to account for this text.
  • Negative sample incorrectly classified as positive: The previous text was classified as having a fairness issue, returning True. However, this item has no issue. Think about why this text was classified as unfair. Improve the original fairness instruction to allow this item by removing or revising restrictions.
This process is run up to $e$ epochs, or stopped early if accuracy reaches 1 or the predictions are stable. We run over $b$ batches of $n$ samples randomly drawn from the training data, using the best-scoring prompt from the final batch for evaluation. [^4] For an example of the process, see Appendix C.
Figure 2: Self-correcting prompt strategy. Data is run through the prompt. If the result is correct, we continue; otherwise, we instruct the LLM to correct the prompt.
@@
We use the validation set to tune the prompts and parameters to optimize the F1 score for each method. Note that for all prompting strategies, the temperature is set to zero; the prompts should only return True or False. Figure 3 shows the best results on the validation set. We explore each model’s effectiveness on unseen data in Section 5.
Figure 3: F1 scores on the validation set for each prompting method. Note that for generic (short)  the F1 score was 0. Full results in Appendix D.
The base generic prompt fails, as the traditional bias and stereotyping issues are less likely to occur in our generated content, and the fairness issues we are concerned with are unlikely to be deemed as problematic out of context. Using a simplified version of our guidelines yields a 0.36 F1 score for identifying fairness issues. The data-driven based on observations in the training data yields much better results (0.70 F1). However, this may not extend well to novel cases, as the prompt is driven purely by our validation data.
Few-shot learning displays some interesting properties: we see significant improvements across all three prompts, using three samples. (This yielded the best results across all validation runs). Even the minimal generic (short) prompt rises to over 0.60 F1 with minimal few-shot prompting.